OpenStudy (anonymous):

The x-intercepts of f(x) = sin (x + π/4) are

OpenStudy (anonymous):

to get x-intercept we take f(x) = 0 sin ( x+ pi/4) = 0 x+ pi/4 =0

OpenStudy (anonymous):

0 and π/4 be the answer

OpenStudy (anonymous):

i think u have to expand that

OpenStudy (anonymous):

i dont think so recheck your answer

OpenStudy (anonymous):

i think it is pie/4

OpenStudy (anonymous):

its two answers because iam working with trig inverses

OpenStudy (anonymous):

\[\it's -\frac{ \Pi }{ 4 }\]

OpenStudy (anonymous):

\[\sin (x + \Pi/4) = \frac{ sinx }{ \sqrt{2} } +\frac{ cosx }{ \sqrt{2} }\]

OpenStudy (anonymous):

\[\frac{ sinx + \cos x }{ \sqrt{2} } = 0\]

OpenStudy (anonymous):

\[sinx + \cos x = 0\] \[\cos^2 x + \sin^2 x = 1\] \[\cos x = \sqrt{1-\sin^2x}\]

OpenStudy (anonymous):

\[sinx = -\sqrt{1-\sin^2x}\] squaring \[\sin^2 x = 1- \sin^2x\] \[2\sin^2 x = 1\] \[sinx = \frac{ 1 }{ \sqrt{2} }\]

OpenStudy (ash2326):

we don't have to complicate this \[f(x)=0=\sin (x+\frac \pi 4)\] so \[x+\frac \pi 4=n\pi\] where n=0,1, 2, 3, 4 so \[x=n\pi-\frac \pi 4\] now we can find the x- intercepts for n=0 \[x=-\frac \pi 4\] for n=1 \[x=\frac {3\pi} 4\] and so on @timtim do you understand this?

OpenStudy (anonymous):

yes i have an idea

OpenStudy (ash2326):

good, now change the value of n and find x intercepts. If you get stuck or get any doubts, do ask me.

OpenStudy (anonymous):

Yup...@ash2326 ...) an Easy one.... from u..)